### v1

@@ Abstract.
The increasing affordability of robot hardware is accelerating the integration of robots into everyday activities. However, training a robot to automate a task typically requires physical robots and expensive demonstration data from trained human annotators. Consequently, only those with access to physical robots produce demonstrations to train robots. To mitigate this issue, we introduce EVE, an iOS app that enables everyday users to train robots using intuitive augmented reality visualizations without needing a physical robot. With EVE, users can collect demonstrations by specifying waypoints with their hands, visually inspecting the environment for obstacles, modifying existing waypoints, and verifying collected trajectories. In a user study ($N=14$, $D=30$) consisting of three common tabletop tasks, EVE outperformed three state-of-the-art interfaces in success rate and was comparable to kinesthetic teaching—physically moving a real robot—in completion time, usability, motion intent communication, enjoyment, and preference ($mean_{p}=0.30$). We conclude by enumerating limitations and design considerations for future AR-based demonstration collection systems for robotics.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; April, 2024; ccs: Human-centered computing Mixed / augmented realityccs: Computer systems organization Robotics
@@ 1. Introduction
The decreasing costs of robot components and systems are accelerating the integration of robots into everyday life. With companies offering robotic arms at a few hundred dollars, portable, lightweight, and mobile robots are being trained to assist individuals with a wide range of everyday tasks (Elephant Robotics, 2023; Robot, 2023; Shaw et al., 2023a). However, training a robot to automate a task often requires expensive manually curated human demonstrations. Annotators are hired and trained to kinesthetically move a robot’s joints and gripper to a sequence of configurations to accomplish a task. For example, to train a robot to clean a table, annotators move the gripper towards a cloth, close the gripper to grasp the cloth, and then begin moving it around to wipe the table.
@@ 4.1. Method
Usability (SUS)
66.3 (20.3) 67.3 (12.3) 70.0 (13.0) 68.3 (15.7)
Motion Intent (rank)
2.9 (1.2) 2.4 (1.0) 1.9 (0.9) 2.8 (1.3)
Enjoyment (rank)
3.5 (1.0)^
2.4 (1.0)
2.0 (0.7)^
2.1 (1.3)
Preference (rank)
3.5 (1.0)^
2.7 (0.8)
1.5 (0.7)^
2.3 (1.1)
@@ 6.2. Findings
Usability (SUS)
86.0 (17.0)
27.3 (17.8)*
59.0 (20.5)*
77.8 (17.0)
Motion Intent (rank)
1.4 (0.7)
3.3 (0.8)*
3.4 (0.7)*
1.9 (0.8)
Enjoyment (rank)
1.5 (0.8)
3.2 (1.0)*
3.3 (0.7)*
2.0 (0.8)
Preference (rank)
1.7 (0.9)
3.4 (0.9)*
3.1 (0.7)*
1.8 (0.7)
@@
Usability (SUS)
83.8 (18.0)
31.3 (21.5)*
66.0 (20.2)*
82.8 (13.9)
Motion Intent (rank)
1.4 (0.7) 3.4 (1.0)
3.2 (0.6)*
2.0 (0.8)
Enjoyment (rank)
2.0 (1.1) 3.0 (1.2)
3.1 (0.9)*
1.9 (1.0)
Preference (rank)
2.2 (1.0) 3.3 (1.1)
2.9 (0.9)*
1.6 (0.8)
@@
Usability (SUS)
88.0 (17.1)*
26.5 (15.5)*
54.0 (20.6)*
71.0 (17.5)
Motion Intent (rank)
1.3 (0.5) 3.1 (0.7)
3.7 (0.5)*
1.9 (0.7)
Enjoyment (rank)
1.2 (0.4)*
3.1 (1.1)
3.5 (0.5)*
2.2 (0.6)
Preference (rank)
1.2 (0.4)
3.5 (0.7)*
3.4 (0.5)*
1.9 (0.6)
@@
Usability (SUS)
86.3 (17.5)
24.3 (17.1)*
56.8 (20.9)*
80.0 (18.7)
Motion Intent (rank)
1.6 (0.8) 3.3 (0.8)
3.3 (0.8)*
1.8 (0.8)
Enjoyment (rank)
1.4 (0.7)
3.5 (0.8)*
3.2 (0.6)*
1.9 (0.7)
Preference (rank)
1.6 (0.8)
3.5 (0.8)*
3.1 (0.7)*
1.8 (0.8)
@@ Appendix A Study Participants Demographics
1 F01 -
Marketing
Somewhat unfamiliar
Somewhat unfamiliar
Chores, remind about deadlines
2 F02 -
Math
Neutral
Somewhat unfamiliar
Chores
3 F03 -
Food Systems
Somewhat unfamiliar
Unfamiliar
Cooking
4 F04 -
Applied Math
Unfamiliar
Unfamiliar
Cooking, driving
5 F05 -
Math
Somewhat unfamiliar
Unfamiliar
Manage files
6 F06 -
Music
Neutral
Unfamiliar
Grab packages, sort folders
7 F07 -
Applied Math
Somewhat familiar
Unfamiliar
Fetch items
8 F08 -
Computer Science
Somewhat unfamiliar
Somewhat unfamiliar
Cooking, driving, grab packages
9 F09 -
Statistics
Somewhat unfamiliar
Unfamiliar
Chores, teaching
10 F10 -
Computer Science
Neutral
Somewhat unfamiliar
Assemble car components
11 - E01
Music
Unfamiliar
Unfamiliar
Grab packages
12 - E02
Electrical Engineering
Somewhat unfamiliar
Unfamiliar
Chores
13 - E03
Statistics
Unfamiliar
Unfamiliar
Fetch items
14 - E04
Business
Somewhat unfamiliar
Somewhat unfamiliar
Chores
15 - E05
Computer Science
Somewhat unfamiliar
Unfamiliar
Fold clothes
16 - E06
Electrical Engineering
Neutral
Somewhat unfamiliar
Cooking
17 - E07
Education
Somewhat unfamiliar
Unfamiliar
Teaching
18 - E08
Informatics
Somewhat unfamiliar
Unfamiliar
Grab packages
19 - E09
Genetics
Neutral
Unfamiliar
Cleaning
20 - E10
Computer Science
Neutral
Neutral
Sort folders
21 - E11
Informatics
Somewhat unfamiliar
Somewhat unfamiliar
Grab packages
22 - E12
Math
Unfamiliar
Unfamiliar
Teaching
23 - E13
Robotics
Familiar
Neutral
Fold clothes
24 - E14
Mechanical Engineering
Familiar
Somewhat familiar
Factory inspection, fetch items

### v2

@@ Abstract.
The increasing affordability of robot hardware is accelerating the integration of robots into everyday activities. However, training robots to automate tasks typically requires physical robots and expensive demonstration data from trained human annotators. Consequently, only those with access to physical robots produce demonstrations to train robots. To mitigate this issue, we introduce EVE, an iOS app that enables everyday users to train robots using intuitive augmented reality visualizations without needing a physical robot. With EVE, users can collect demonstrations by specifying waypoints with their hands, visually inspecting the environment for obstacles, modifying existing waypoints, and verifying collected trajectories. In a user study ($N=14$, $D=30$) consisting of three common tabletop tasks, EVE outperformed three state-of-the-art interfaces in success rate and was comparable to kinesthetic teaching—physically moving a real robot—in completion time, usability, motion intent communication, enjoyment, and preference ($mean_{p}=0.30$). We conclude by enumerating limitations and design considerations for future AR-based demonstration collection systems for robotics.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; , 2024; ccs: Human-centered computing Mixed / augmented realityccs: Computer systems organization Robotics
@@ 1. Introduction
The decreasing costs of robot components and systems are accelerating the integration of robots into everyday life. With companies offering robotic arms at a few hundred dollars, portable, lightweight, and mobile robots are being trained to assist individuals with a wide range of everyday tasks (Elephant Robotics, 2023; Robot, 2023; Shaw et al., 2023a). However, training robots to automate tasks often requires expensive manually curated human demonstrations. Annotators are hired and trained to kinesthetically move a robot’s joints and gripper to a sequence of configurations to accomplish a task. For example, to train a robot to clean a table, annotators move the gripper towards a cloth, close the gripper to grasp the cloth, and then begin moving it around to wipe the table.
@@ 4.1. Method
Usability (SUS) 66.3 (20.3) 67.3 (12.3) 70.0 (13.0) 68.3 (15.7)
Motion Intent (rank) 2.9 (1.2) 2.4 (1.0) 1.9 (0.9) 2.8 (1.3)
Enjoyment (rank) 3.5 (1.0)^ 2.4 (1.0) 2.0 (0.7)^ 2.1 (1.3)
Preference (rank) 3.5 (1.0)^ 2.7 (0.8) 1.5 (0.7)^ 2.3 (1.1)
@@ 6.2. Findings
Usability (SUS) 86.0 (17.0) 27.3 (17.8)* 59.0 (20.5)* 77.8 (17.0)
Motion Intent (rank) 1.4 (0.7) 3.3 (0.8)* 3.4 (0.7)* 1.9 (0.8)
Enjoyment (rank) 1.5 (0.8) 3.2 (1.0)* 3.3 (0.7)* 2.0 (0.8)
Preference (rank) 1.7 (0.9) 3.4 (0.9)* 3.1 (0.7)* 1.8 (0.7)
@@
Usability (SUS) 83.8 (18.0) 31.3 (21.5)* 66.0 (20.2)* 82.8 (13.9)
Motion Intent (rank) 1.4 (0.7) 3.4 (1.0) 3.2 (0.6)* 2.0 (0.8)
Enjoyment (rank) 2.0 (1.1) 3.0 (1.2) 3.1 (0.9)* 1.9 (1.0)
Preference (rank) 2.2 (1.0) 3.3 (1.1) 2.9 (0.9)* 1.6 (0.8)
@@
Usability (SUS) 88.0 (17.1)* 26.5 (15.5)* 54.0 (20.6)* 71.0 (17.5)
Motion Intent (rank) 1.3 (0.5) 3.1 (0.7) 3.7 (0.5)* 1.9 (0.7)
Enjoyment (rank) 1.2 (0.4)* 3.1 (1.1) 3.5 (0.5)* 2.2 (0.6)
Preference (rank) 1.2 (0.4) 3.5 (0.7)* 3.4 (0.5)* 1.9 (0.6)
@@
Usability (SUS) 86.3 (17.5) 24.3 (17.1)* 56.8 (20.9)* 80.0 (18.7)
Motion Intent (rank) 1.6 (0.8) 3.3 (0.8) 3.3 (0.8)* 1.8 (0.8)
Enjoyment (rank) 1.4 (0.7) 3.5 (0.8)* 3.2 (0.6)* 1.9 (0.7)
Preference (rank) 1.6 (0.8) 3.5 (0.8)* 3.1 (0.7)* 1.8 (0.8)
@@ Appendix A Study Participants Demographics
1 F01 - Marketing Somewhat unfamiliar Somewhat unfamiliar Chores, remind about deadlines
2 F02 - Math Neutral Somewhat unfamiliar Chores
3 F03 - Food Systems Somewhat unfamiliar Unfamiliar Cooking
4 F04 - Applied Math Unfamiliar Unfamiliar Cooking, driving
5 F05 - Math Somewhat unfamiliar Unfamiliar Manage files
6 F06 - Music Neutral Unfamiliar Grab packages, sort folders
7 F07 - Applied Math Somewhat familiar Unfamiliar Fetch items
8 F08 - Computer Science Somewhat unfamiliar Somewhat unfamiliar Cooking, driving, grab packages
9 F09 - Statistics Somewhat unfamiliar Unfamiliar Chores, teaching
10 F10 - Computer Science Neutral Somewhat unfamiliar Assemble car components
11 - E01 Music Unfamiliar Unfamiliar Grab packages
12 - E02 Electrical Engineering Somewhat unfamiliar Unfamiliar Chores
13 - E03 Statistics Unfamiliar Unfamiliar Fetch items
14 - E04 Business Somewhat unfamiliar Somewhat unfamiliar Chores
15 - E05 Computer Science Somewhat unfamiliar Unfamiliar Fold clothes
16 - E06 Electrical Engineering Neutral Somewhat unfamiliar Cooking
17 - E07 Education Somewhat unfamiliar Unfamiliar Teaching
18 - E08 Informatics Somewhat unfamiliar Unfamiliar Grab packages
19 - E09 Genetics Neutral Unfamiliar Cleaning
20 - E10 Computer Science Neutral Neutral Sort folders
21 - E11 Informatics Somewhat unfamiliar Somewhat unfamiliar Grab packages
22 - E12 Math Unfamiliar Unfamiliar Teaching
23 - E13 Robotics Familiar Neutral Fold clothes
24 - E14 Mechanical Engineering Familiar Somewhat familiar Factory inspection, fetch items